OpenStudy (anonymous):

A golf ball is hit from the ground with an initial speed of 192 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground? 39 secs 19 secs 12 secs 6 secs

OpenStudy (anonymous):

assume g = -32 , u = 192, s = 0, t = ? using the formula s = ut + 0.5gt^2 0 = 192t - 16 t^2 16t( t - 12) = 0 t = 12 seconds

OpenStudy (anonymous):

Tysm

OpenStudy (anonymous):

If you can help me with the rest of these Id be so grateful

OpenStudy (anonymous):

i haven't a lot of time but give me one more

OpenStudy (anonymous):

A juggler is performing an act by juggling several balls. The juggler throws the balls up at an initial height of 3.9 feet with a speed of 14.8 feet per second. If the juggler did not catch a ball, about how long will it take the ball to hit the floor? 1.45 secs 1.14 secs 0.28 secs 0.21 secs

OpenStudy (anonymous):

again g = -32 ft per sec, s = displacement = -3.9, u = 14.8, t = ? -3.9 = 14.8t - .5 * 32 * t^2 16t^2 - 14.8t - 3.9 = 0 t = -o.21 or 1.14 time cant be negative so t must be 1.14 secs

OpenStudy (anonymous):

Thats what I thought it was cool

OpenStudy (anonymous):

yeah you need the constant acceleration equations to solve these type problems they are v = u + at v = final velocity, u = initial , a = acceleration ( g= accn of gravity) s = ut + 0.5at^2 v^2 = u^2 + 2as s =[ (u+v)t] / 2 s = displacement